OpenStudy (anonymous):

Please help with this problem!! Given - f '' (x) = x^3 + 10x^2 + 6x f ' (x) = -4 f (x) = -10 Find f ' (x) = _________ f (x) = __________

OpenStudy (gorv):

on integrating f ''( x) f ' (x)=x^4/4 +10x^3/3+6x^2/2 on integrating this f(x)= x^5/(4*5) + (10/3)*x^4/4+3x^3/3 f(x)=(x^5)/20 +(10/12)*x^4 +x^3 f(x)=(x^5)/20 +(5/6)*x^4 +x^3

OpenStudy (gorv):

@MG757983